OpenStudy (bossimbacon):

Which answer shows this equation written in standard form? 4 – 3(x – y) = –8x + 2 A. –11x + 3y = –2 B. 5x – 3y = 2 C. 5x + 3y = –2 D. 5x + y = 2

OpenStudy (haseeb96):

ax+by+c=0 this is the example of standard form so when we solve this above question 4-3(x-y)=-8x+2 -3(x-y) =-8x+2 -4 now multiply the left hand side constant ''3'' with x and y it would be -3x +3y =-8x -2 now bring the -8x and -2 to the left hand side and change the sign also of it -3x +8x +3y +2 =0 5x +3y +2 =0 this is your standard form u can see this standard form which is similar to my example equation

OpenStudy (haseeb96):

now according to your option u can write your answer in this form 5x + 3y =-2 okay for question's answer